help solve 5 3/4 + 2 2/3 =
help solve 2 3/10 + 3 4/5 =
help solve 3 3/4 + 1 1/2 =

Found 2 solutions by kekegeter , Alan3354:
Answer by kekegeter (97)

(5*4)+3/4 +(2*3)+2/3
101/2^2*3
101/12

Answer by Alan3354(69443)

help solve 5 3/4 + 2 2/3
This is not solving, it's adding.
5 3/4 + 2 2/3
Get rid of mixed numbers, they're useless.
= 23/4 + 8/3
Find a common denominator. The product always works. 4*3 = 12
23/4 + 8/3 = 69/12 + 32/12 = (69+32)/12
= 101/12
----------------------
help solve 2 3/10 + 3 4/5
Do it the same as the 1st one.
email via the TY note for help or to check your work.